GRADUATE STUDENT ASSIGNMENTS (Finalizes assignments for last four weeks)

June 17 – 28  July 1 – 12  July 15 – 26  July 29 – Aug 9
Approximations  Sally  Sally
Discrete  Babai  Babai
Geo  Farb  May  Rothenberg  Brock
Quantum mechanics  Constantin  and Kiselev
Invariant Theory  Kottwitz
Categorification  May
Shreya Amin  —————-  —————-  QM  QM
Angela Barnhill  Geo  Geo ———-  ———- Geo  ———————-
David Balduzzi  Sally  Sally  Kottwitz  May
Nils Barth  Babai  Babai  ———————- ———————-
Mark Behrens  Geo  Geo  Geo  Geo
Ben Blander  Geo  Geo  Geo  Geo
Andrew Blumberg  Geo  Geo  Geo  Geo
Chris Bremer  Babai  ———————-  Kottwitz  May
Kariane Calta  ———————- ———————- ———————-  Geo
Jeremy Copeland  ———————- ———————-  QM  QM
Gautam Iyer  Sally  Sally  QM  ———————-
Nick Gursky  ———————-  Geo  Kottwitz  May
Sonal Jain  Babai  ———————– ———- Kottwitz  May
Joe Johns  ———————- ———————-  QM  QM
Ben Lee  Geo  Geo  Geo  Geo
Vincent Lucarelli  Sally  ———————- ———————- ———————-
Joshua Maher  Geo  Geo  Geo  Geo
Dan Margalit  Geo  Geo  ———————-  Geo
Karin Melnick  Babai  Babai  ———————- ———————-
Nick Michalowski  Babai  Babai  QM  QM
Deepam Patel  Sally  Sally  Kottwitz  May
Rochelle Pereira  ———————-  Geo  Geo  ———————-
Ann Scheels  ———————- ———————-  QM  QM
Steve Spallone  ———- Sally  Sally  Kottwitz  ———————-
Loren Spice  Sally  Sally  ———————- ———————-